OpenStudy (anonymous):

PLease help! What are the foci of the ellipse? Graph the ellipse. x^2/49+y^2/64=1

OpenStudy (amistre64):

well, textbooks like to use "c" for focal length, ill use "f" b^2 + f^2 = a^2 therefore, f^2 = a^2 - b^2, sqrt for f the bottoms of your fractions there are the a^2 and b^2 parts. a^2 > b^2

OpenStudy (amistre64):

since your y parts are bigger then your x parts ... your focuses will be along the y axis (0,f) and (0,-f) in this case since its also centered at the origin